# EEG Foundation Models: A Critical Review of Current Progress and Future Directions

### v2

@@ Abstract
Patterns of electrical brain activity recorded via electroencephalography (EEG) offer immense value for scientific and clinical investigations. The inability of supervised EEG encoders to learn robust EEG patterns and their over-reliance on expensive signal annotations have sparked a transition towards general-purpose self-supervised EEG encoders, i.e., EEG foundation models (EEG-FMs), for robust and scalable EEG feature extraction. However, the real-world readiness of early EEG-FMs and the rubric for long-term research progress remain unclear. A systematic and comprehensive review of first-generation EEG-FMs is therefore necessary to understand the current state-of-the-art and identify key directions for future EEG-FMs. To that end, this study reviews 10 early EEG-FMs and presents a critical synthesis of their methodology, empirical findings, and outstanding research gaps. We find that most EEG-FMs adopt a sequence-based modeling scheme that relies on transformer-based backbones and the reconstruction of masked sequences for self-supervision. However, model evaluations remain heterogeneous and largely limited, making it challenging to assess their practical off-the-shelf utility. In addition to adopting standardized and realistic evaluations, future work should demonstrate more substantial scaling effects and make principled and trustworthy choices throughout the EEG representation learning pipeline. We believe that developing benchmarks, software tools, technical methodologies, and applications in collaboration with domain experts may further advance the translational utility and real-world adoption of EEG-FMs.
@@ 1 Introduction
Electroencephalography (EEG) is a versatile tool for recording the brain’s electrical activity and offers immense value to neuroscientists, biomedical engineers, and clinicians [1]. The interpretation of multi-channel EEG recordings has historically been performed by trained experts because of their rich information content in the spatial and temporal dimensions. Despite decades of research into quantitative EEG features derived based on domain expertise, expert visual interpretation remains the gold standard for clinical EEG evaluation even to this day[2]. However, in the last decade, EEG research has seen significant advances in deep learning-based approaches for extracting application-specific features from raw EEG data (EEG-DL) [3, 4]. Although this line of research held significant potential to augment traditional visual EEG review, it met with limited success due to several reasons. Much of the EEG-DL research relied on supervised learning on a very narrow group of EEG tasks that resulted from laborious expert-driven annotation efforts [5]. These efforts were clearly unscalable to cover the wide range of EEG tasks that might be needed in clinical or scientific EEG review. Furthermore, such supervised EEG encoders were highly susceptible to overfitting erroneous and noisy instances in training data, causing concerns regarding robustness and transferability to other tasks and datasets. This lack of robustness was further exacerbated by the variability of EEGs across recording sites, acquisition systems, subjects, and sessions, which led to an overall lack of trust in supervised EEG encoders [6, 7]. These limitations emphasized a need for EEG-DL models that rely less on expert EEG labels and yield robust and trustworthy EEG-DL models with high translational value.
@@
The emerging paradigm of foundation models (FMs)[8], based on label-free self-supervised learning (SSL) and efficient transfer learning, is a promising solution for these data-related challenges. Similar to the mainstream vision [9, 10, 11] and language [12, 13, 14] FMs, EEG foundation models (EEG-FMs) are trained to identify salient EEG features from raw unlabeled EEG recordings by using various SSL pretext tasks, such as masked reconstruction or masked prediction, via a process known as pretraining. EEG-FMs learn to represent EEG data as compressed embeddings in a latent space by leveraging the intrinsic properties found in the raw data. Pretrained EEG-FMs can then be adapted for various downstream applications using only very small amounts of labeled data, thereby alleviating the burden of expert EEG annotations – the main bottleneck in supervised EEG-DL research. As such, EEG-FMs hold promise as powerful, off-the-shelf feature extractors (or encoders) that can support scientific research, next-generation robust brain-computer interfaces, and augmented neurological decision support.
@@
Several first-generation EEG-FMs have been proposed over the last few years [15, 16, 17, 18, 19, 20, 21, 22, 23, 24], whose cumulative count is shown in Figure 1. Despite the growing interest, many questions remain unanswered regarding the design choices within EEG-FMs, the learned representations, the performance on various real-world applications, and the overall guarantees on robustness and trustworthiness. For example, the choice of EEG input representations, the architectural components, and the SSL pretext tasks can vary significantly between models, and the effects of those choices on the learned features are unclear. The complexity, quality, and flexibility of the representations learned by EEG-FMs and their relation to brain physiology have not been sufficiently studied. Furthermore, the performance and generalizability of these EEG-FMs beyond the common public datasets and benchmarks have not been adequately evaluated. These questions and concerns call for a comprehensive review of the first-generation EEG-FMs focusing on the various architectural choices, pretraining approaches, evaluations, and trustworthiness aspects, to identify the rubrics for meaningful long-term progress in EEG-FM research and advance their translational value.
Figure 1: Cumulative EEG-FM search results from 2021 to September 30th, 2024. EEG-FMs are displayed chronologically in the plot based on their preprint publication dates.
@@
To that end, this topical review focuses on the first generation of EEG-FMs and critically analyzes their building blocks, identifies key takeaways and research gaps, and suggests directions for future EEG-FM research. The contributions of this review are two-fold: a) it offers a principled understanding of EEG-FMs and the current state-of-the-art in a manner accessible to both EEG scientists and engineers, and b) it provides a domain-centric perspective and critique of existing work that can guide future generations of EEG-FMs. This review is organized as follows. Section 2 introduces the EEG recording procedures, the information content of EEGs, traditional quantitative EEG approaches based on feature engineering, the sources of data variability, and the different functional components of foundation models. Section 3 describes the search strategy we undertook to identify previously published EEG-FMs and provides brief summaries of those models. Section 4 provides a comprehensive analysis of each of the identified EEG-FMs along several dimensions, including training data, data preprocessing, input representations, model architectural components, and model evaluation strategies. Section 5 discusses the key takeaways and research gaps, and Section 6 lays out our view for future directions that would further advance the current state of the art.
@@ 2 Background
EEG is a common neurophysiological technique used to record electrical activity in the brain. It is widely used in clinical and research settings for diagnosing neurological disorders[25], studying brain function[26], and developing brain-computer interfaces[27]. EEG measures voltage fluctuations resulting from ionic current flow within neurons[28] and is typically captured using electrodes placed directly on the scalp or intracranially. While intracranial EEG (iEEG) measures summated action potentials generated by small neuronal populations, scalp EEG measures the electrical activity generated by large neuronal populations that pass through multiple layers of the brain and head (i.e., cerebrospinal fluid, meninges, skull, scalp). The conduction of neuronal activity through these layers causes significant attenuation and spatial mixing of the EEG signal[29]. As a result, scalp EEG has considerably different spatial resolution, amplitude, frequency, and noise content compared to intracranial EEG[30].
@@ 2.1 Contents and Sources of Variability in EEG Data
The EEG signal is a noisy, high-dimensional, spatio-temporal measurement of brain electrical activity. Several factors influence the contents of EEG recordings, including physiological, pathological, and artifactual components. Physiological elements: The physiological elements present in EEG can be described using the various oscillatory components (delta, theta, alpha, beta, and gamma bands)[31]. The general attentional state of the subject during the recording (awake, drowsy, asleep) influences the frequency characteristics of EEG[32]. In addition, EEG also shows changes related to natural aging[33] and other benign changes in brain structure and function.
Pathological elements: The primary diagnostic value of EEG is due to its ability to capture various pathological brain activity patterns. These electrographic patterns include seizures, interictal epileptiform discharges (e.g., spikes and sharp waves), triphasic waves, rhythmic discharges, lateralized periodic discharges, background slowing, and other diffuse or focal abnormal electroencephalogram patterns[25].
Artifactual elements: EEG artifacts can be introduced by biological phenomena (e.g., eye blinks, head movement, muscle activity, cardiac signals), recording conditions (e.g., signal discontinuities, transient filter effects), or external sources (e.g., electromagnetic interference)[34]. Eye- and muscle-related activities are the primary sources of artifacts found in EEGs, which introduce slow, high-amplitude, and fast, low-amplitude artifactual changes, respectively. In addition to these artifacts, certain medications, such as anti-depressants and anti-seizure medications, affect the EEG as well. For example, benzodiazepines and barbiturates are known to cause changes to the EEG[35].
The various practices and configurations used in clinical EEG studies introduce variabilities across EEG datasets. First, the hardware used to record EEGs can have preset configurations for filtering, analog-to-digital quantization, amplification magnitude, and broadband noise levels[36]. Second, the EEG examination protocol may differ between clinical sites: the recording length could be different; recordings can have variable amounts of wake/sleep segments and eyes open/closed segments; and they may include cognitive tests and activation procedures such as sleep deprivation, photic stimulation, or hyperventilation. In addition, the EEG layout (i.e., number of leads and their positions) and reference leads may vary among sites. Finally, anatomical[37], genetic[38], and biochemical[39] differences introduce inter-individual variability. These factors make data ingestion and standardization for group-level modeling non-trivial, especially in intracranial EEG datasets, where lead locations are subject-specific.
@@ 2.2 Feature Extraction from EEG Signals
Characteristics of clinical EEGs are assessed visually by experts trained to detect pathological events and adverse deviations from healthy brain activity[40]. However, in non-clinical settings, the EEG is characterized by signal features derived computationally and analytically from the raw data. Quantitative EEG features serve as lower-dimensional proxies of the raw signal that quantify pertinent brain activity and provide objective EEG assessments for various applications. Classical EEG features are derived from signal processing, statistics, and information theory domains. Examples include relative band-limited power, power ratios, independent components, sample entropy, and kurtosis, among others[41]. In select instances, classical EEG features isolate and quantify meaningful characteristics of underlying brain physiology, such as neuronal oscillations and functional connectivity, and are therefore considered robust and interpretable. In supervised deep learning (DL) models, EEG features are learned directly from raw EEG data to increase performance for a particular task[42]. While EEG-DL models may push the boundaries of task performance, their black-box nature makes it difficult, or at least non-trivial, to understand their meaning and relevance to brain physiology. Furthermore, EEG-DL features may overfit the noise in EEG signals, making them less likely to generalize out of the box. Hence, the robustness, reliability, and interpretability of EEG-DL feature extractors remain open research problems.
@@ 2.3 Pillars of Foundation Modeling
In this section, we deconstruct the FM monolith into a set of universal, domain-agnostic building blocks/pillars. This set is minimally required for the systematic comparison of diverse existing EEG-FMs and to build new ones.
@@
Representation of input data: EEG signals are acquired as a spatio-temporal data stream. Visual review of clinical EEGs is conducted primarily in this ’native’ view. However, computational interpretation/analysis of EEGs can benefit from alternative, perhaps more informative, representations of raw data. For example, representing EEGs as a temporal stream of short spatio-spectral data segments (10s each) could make it easier for EEG-FMs to learn latent frequency patterns. Alternatively, representing EEGs using spatial covariance matrices can make it easier to extract latent connectivity patterns. Broadly, the choice/format of data representation at the input stage may greatly influence the quality and semantics of learned latent representations[43]. In biosignal domains where data are sampled from underlying physiological processes, constructing informative input data views may itself require significant offline transformations or even domain-specific modeling, such as inverse or source modeling[44]. Without these offline efforts, data-driven learning of latent representations via FMs may be inefficient or suboptimal.
@@
Architectural and functional design of FMs: At an architectural level, FMs are typically built from several non-linear transformer blocks that can be readily stacked to scale the model depth and size. Transformer blocks themselves contain multiple feed-forward layers followed by a multi-head attention module[45]. Vision FMs may use convolutional blocks to exploit the desirable spatial inductive bias of convolutions[46]. Regardless of model size and any low-level architectural choices made, encoder-only FMs are understood to have only two high-level functional components: a backbone network and a final single fully connected layer. The backbone network serves as a feature extractor, providing low-dimensional embeddings (or representations) of the input data.The final layer, referred to as a task head, then uses these embeddings for task-specific purposes such as classification or regression.
@@
Self-supervised learning (SSL): In the SSL paradigm, the supervision for training is strategically derived from the unlabeled input data (pseudo-labels) rather than from human-curated labels as is done in supervised learning[47]. SSL objectives (pretext tasks) are commonly either contrastive or generative in nature. Contrastive pretext tasks use meaningfully augmented or noisy views of the input to learn robust data representations using encoder-only backbones[48]. Generative pretext tasks mask various portions of the input and reconstruct/generate the masked portions using encoder and decoder-based backbones[49]. In either case, SSL learns general intrinsic relationships within the data that help create a task-agnostic backbone network that serves as a general feature extractor. The structure of representations learned by SSL tasks and general principles for effective SSL task design are under active study. Regardless, the value of SSL is particularly high in the EEG domain as it improves model generalizability[50] and partially alleviates the need for expert EEG annotations that are expensive to obtain, subjective, and error-prone[7].
@@
Transfer learning and model adaptation: The backbone network, whether trained with or without self-supervision, allows for the re-use, transfer, and adaptation of EEG features (i.e., data-driven EEG knowledge) across semantically overlapping tasks[51]. The proportion of the backbone adapted for a specific application is flexible. For example, the backbone may be entirely fixed/frozen (linear probing), partially updated, or entirely updated (fine-tuning)[52]. This adaptation process may utilize slower learning rates or layer-specific learning rates to maximally preserve pretrained knowledge within the backbone. In either case, any pretext task-specific layers at the top are discarded and replaced with a new trainable linear head that performs classification or regression for the ‘downstream’ application. Using application-specific data splits and human-curated task-specific labels, the full model is then trained further until convergence. The development of algorithms for robust and efficient adaptation is an active area of research. Broadly, transfer learning and model adaptation techniques obviate the need to train new EEG-DL models from scratch by leveraging knowledge from pre-existing models.
@@
Data and model scale: The scaling up of pretraining data and model size has contributed towards improved pretrained latent features, increased downstream task performance, sample efficiency, and model generalizability, among other properties[53, 54, 55, 56, 57, 58]. Empirical scaling laws have established power-law relationships between FM error rates and three key modeling levers, namely the pretraining data size, model size, and compute[55, 59], which suggest that an exponential increase in data and model size can indeed improve FM performance, albeit with diminishing returns. It is noteworthy that vision and language FMs are often multi-billion-parameter models and are pretrained on internet-scale datasets[60, 61, 62]. Overall, the balance between data and model size to maximize performance under a fixed compute budget remains a critical consideration in FM research.

### v3

@@ Abstract
Premise. Patterns of electrical brain activity recorded via electroencephalography (EEG) offer immense value for scientific and clinical investigations. The inability of supervised EEG encoders to learn robust EEG patterns and their over-reliance on expensive signal annotations have sparked a transition towards general-purpose self-supervised EEG encoders, i.e., EEG foundation models (EEG-FMs), for robust and scalable EEG feature extraction. However, the real-world readiness of early EEG-FMs and the rubrics for long-term research progress remain unclear. Objective. In this work, we conduct a review of ten early EEG-FMs to capture common trends and identify key directions for future development of EEG-FMs. Methods. We comparatively analyze each EEG-FM using three fundamental pillars of foundation modeling, namely the representation of input data, self-supervised modeling, and the evaluation strategy. Based on this analysis, we present a critical synthesis of EEG-FM methodology, empirical findings, and outstanding research gaps. Results. We find that most EEG-FMs adopt a sequence-based modeling scheme that relies on transformer-based backbones and the reconstruction of masked temporal EEG sequences for self-supervision. However, model evaluations remain heterogeneous and largely limited, making it challenging to assess their practical off-the-shelf utility. In addition to adopting standardized and realistic evaluations, future work should demonstrate more substantial scaling effects and make principled and trustworthy choices throughout the EEG representation learning pipeline. Significance. Our review indicates that the development of benchmarks, software tools, technical methodologies, and applications in collaboration with domain experts may advance the translational utility and real-world adoption of EEG-FMs.
@@ 1 Introduction
Electroencephalography (EEG) is a widely used neurophysiological modality for measuring the brain’s electrical activity with high temporal resolution [mushtaqOneHundredYears2024]. Multi-channel EEG recordings encode complex neural dynamics across spatial, temporal, and spectral dimensions, necessitating interpretation by highly trained experts. Despite extensive efforts to develop quantitative EEG representations using handcrafted features, advanced signal processing, and contemporary machine learning approaches, expert visual review remains the gold standard for clinical EEG interpretation and decision-making. However, in the last decade, EEG research has seen significant advances in deep learning-based approaches for extracting application-specific features from raw EEG data (EEG-DL) [royDeepLearningbasedElectroencephalography2019a, craikDeepLearningElectroencephalogram2019]. Although this line of research held significant potential to augment traditional visual EEG review, it met with limited success due to several reasons. Much of the EEG-DL research comprised end-to-end learning methods that relied on supervised learning and were trained on a very narrow group of EEG tasks, i.e., use-cases, and datasets that resulted from laborious expert-driven annotation efforts [obeid2016temple]. These efforts were clearly unscalable to cover the wide range of EEG tasks that might be needed in clinical or scientific EEG review. Furthermore, such supervised EEG encoders were highly susceptible to overfitting on erroneous and noisy training instances, raising concerns about robustness and transferability to other tasks and datasets. This lack of robustness was further exacerbated by the variability of EEG recordings across recording sites, acquisition systems, subjects, and sessions, leading to a lack of trust in supervised EEG encoders [wagh2022evaluating, banville2021uncovering]. These limitations emphasized the need for EEG-DL solutions that rely less on expert EEG labels and yield robust, trustworthy, and explainable models with high translational value.
@@
The emerging paradigm of foundation models (FMs) [bommasani2021opportunities], based on label-free self-supervised learning (SSL) and efficient transfer learning, is a promising solution for these data-related challenges. Similar to the mainstream vision [yuan2021florence, kirillov2023segment, radford2021learning] and language [chowdhery2023palm, touvron2023llama, achiam2023gpt] FMs, EEG foundation models (EEG-FMs) are trained to identify salient EEG features from raw unlabeled EEG recordings by using various SSL pretext tasks, such as masked reconstruction or masked prediction, via a process known as pretraining. EEG-FMs learn to represent EEG data as compressed embeddings in a latent space by leveraging the intrinsic properties found in the raw data. Pretrained EEG-FMs can then be adapted for various downstream applications using only very small amounts of labeled data, thereby alleviating the burden of expert EEG annotations – the main bottleneck in supervised EEG-DL research. As such, EEG-FMs hold promise as powerful, off-the-shelf feature extractors (or encoders) that can support scientific research, next-generation robust brain-computer interfaces, and augmented neurological decision support.
@@
Several EEG-FMs have been proposed over the last few years [wang2023brainbert, cui2024neuro, zhang2024brant, yang2024biot, chen2024eegformer, jiang_large_2024, panchavati2024mentality, jiang_neurolm_2024, shi_fome_2024, yuan2024brainwavebrainsignalfoundation], whose chronological order along with total annual Google Scholar search results for the term “EEG Foundation Model” is shown in Figure 1. Despite the growing interest, many questions remain unanswered regarding the design choices in EEG-FMs, the learned representations, performance across various real-world applications, and the guarantees of robustness and trustworthiness. For example, the choice of EEG input representations, architectural components, and SSL pretext tasks can vary significantly across models, and the effects of those choices on the learned features are unclear. The complexity, quality, and flexibility of the representations learned by EEG-FMs and their relation to brain physiology have not been sufficiently studied, raising questions about their explainability. Furthermore, the performance and generalizability of these EEG-FMs beyond the common public datasets and benchmarks have not been adequately evaluated. These questions and concerns call for a critical review of the early EEG-FMs, focusing on architectural choices, pretraining approaches, evaluations, and trustworthiness, to identify rubrics for meaningful long-term progress in EEG-FM research and to advance their translational value.
Figure 1: Electroencephalography foundation model (EEG-FM) publication trends. The total annual Google Scholar search results for the term “EEG Foundation Model” between 2021 and 2024 (September) are shown in blue. The specific EEG-FMs reviewed in this study (i.e., according to the search criteria described in section 3) are listed in red, in chronological order by their preprint publication dates.
@@
To that end, this topical review surveys ten early EEG-FMs (between January 2021 and September 2024) and critically analyzes their building blocks, identifies key takeaways and research gaps, and suggests directions for future EEG-FM research. Recent studies have begun to survey EEG-FMs [yuxuan2025foundation, zhou2025brain, weng2025self], but these reviews primarily emphasize technical components such as model architectures and self-supervised learning strategies. In contrast, critical factors related to data representation, evaluation scope, and rigor have received relatively little attention. Even more limited is the discussion of real-world translational requirements and the practical constraints under which EEG-FMs will ultimately need to operate. In this review, we offer a more holistic perspective on the development and assessment of EEG-FMs, adopting a data- and domain-centric viewpoint to evaluate their maturity for clinical translation and real-world use. We also highlight complementary research directions that can help position future EEG-FMs for meaningful practical impact. This review is organized as follows. Section 2 introduces the EEG recording procedures, the information content of EEGs, traditional quantitative EEG approaches based on feature engineering, the sources of data variability, and the different functional components of foundation models. Section 3 describes the search strategy we undertook to identify previously published EEG-FMs and provides brief summaries of those models. Section 4 provides a comprehensive analysis of each of the identified EEG-FMs along several dimensions, including training data, data preprocessing, input representations, model architectural components, and model evaluation strategies. Section 5 discusses the key takeaways and research gaps, and Section 6 lays out our view for future directions that would further advance the current state of the art.
@@ 2 Background
EEG is a common neurophysiological technique for recording electrical activity generated by the brain using one or more electrodes. EEG emerged in the early nineteenth century and has since expanded into many areas of science and medicine. It is widely used in clinical and research settings for diagnosing neurological disorders [tatum2018clinical], studying brain function [da2013eeg], and developing brain-computer interfaces [varbu2022past]. EEG measures voltage potentials and their fluctuations resulting from ionic current flow mainly generated by firing populations of neurons [buzsaki2012origin]. Scalp EEG is typically captured using electrodes placed directly on the scalp. EEG can also be recorded invasively, with electrodes placed intracranially (within the brain) or subdurally (under the skin), commonly referred to as intracranial EEG (iEEG). While iEEG measures summated local field potentials generated by smaller neuronal populations, scalp EEG measures the electrical activity generated by large neuronal populations that pass through multiple layers of the brain and head (i.e., cerebrospinal fluid, meninges, skull, muscles, scalp). The conduction of neuronal activity through these layers causes significant attenuation and spatial mixing of the EEG signal [brunner2016volume], including artifacts caused by movements of electrodes and muscle activity. As a result, scalp EEG has considerably different spatio-temporal resolution, amplitude, frequency, and noise content compared to intracranial EEG [petroff2016comparison].
@@ 2.1 Contents and Sources of Variability in EEG Data
The EEG signal is a noisy, high-dimensional, spatio-temporal measurement of electrical brain activity. Several factors influence the contents of EEG recordings, including physiological, pathological, and artifactual elements. Physiological elements: Most of the physiological elements present in EEG can be described using the various oscillatory components known as the delta, theta, alpha, beta, and gamma bands, each with distinct oscillatory frequencies and physiological mechanisms [klimesch1996memory]. The general attentional state of the subject during the recording (awake, drowsy, asleep) influences the spatio-temporal characteristics of EEG [oken2006vigilance]. In addition, EEG also shows changes related to natural aging [perinelli2022power] and other benign changes in brain structure and function.
Pathological elements: The primary diagnostic value of EEG is due to its ability to capture various pathological brain activity patterns. These electrographic patterns include seizures, interictal epileptiform discharges (e.g., spikes and sharp waves), triphasic waves, rhythmic discharges, lateralized periodic discharges, background slowing, and other diffuse or focal abnormal electroencephalogram patterns [tatum2018clinical].
Artifactual and noise elements: EEG artifacts can be introduced by biological phenomena (e.g., eye blinks, head movement, muscle activity, cardiac signals), recording conditions (e.g., signal discontinuities, transient filter effects), or external sources (e.g., electromagnetic interference) [amin2023normal]. Eye- and muscle-related activities are the primary sources of artifacts in EEGs, introducing slow, high-amplitude changes and fast, low-amplitude changes, respectively. In addition to these artifacts, consistent noise from the power line and its harmonics may also be present. Finally, certain medications, such as anti-depressants and anti-seizure medications, affect the EEG as well. For example, benzodiazepines and barbiturates are known to cause changes to the EEG [blume2006drug].
The various practices and configurations used in clinical studies introduce variabilities across EEG datasets. First, the hardware used to record EEGs can have preset configurations for sampling frequency, filtering, analog-to-digital quantization, amplification magnitude, and broadband noise levels [li2021review]. Second, the EEG examination protocol may differ between clinical sites: the recording length could be different; recordings can have variable amounts of wake/sleep segments and eyes open/closed segments; and they may include cognitive tests and activation procedures such as sleep deprivation, photic stimulation, or hyperventilation. In addition, the EEG layout (i.e., the number of leads and their positions) and the reference leads may vary across sites. Finally, anatomical [cespedes2020influence], genetic [smit2012individual], and biochemical [gao2017inferring] differences introduce inter-individual variability. These factors make data ingestion and standardization non-trivial, especially in intracranial EEG datasets, where lead locations are subject-specific.
@@ 2.2 Feature Extraction from EEG Signals
Characteristics of clinical EEGs are assessed visually by experts trained to detect pathological events and adverse deviations from healthy brain activity [tatum2021handbook]. However, in non-clinical settings, the EEG is characterized by signal features derived computationally or analytically from the raw data. Quantitative EEG features serve as lower-dimensional proxies of the raw signal, quantifying pertinent brain activity and providing objective EEG assessments for various applications. Classical EEG features are derived from signal processing, statistics, and information theory domains. Examples include relative band-limited power, power ratios, independent components, sample entropy, and kurtosis, among others [popa2020role]. In select instances, these EEG features isolate and quantify meaningful characteristics of underlying brain physiology, such as neuronal oscillations and functional connectivity, and are therefore considered robust and interpretable. In supervised EEG-DL models, EEG features are learned directly from raw EEG data to increase performance for a particular task [roy2019deep]. While EEG-DL models may push the boundaries of task performance, their black-box nature makes it non-trivial to relate them to brain physiology. Furthermore, EEG-DL features may overfit the noise in EEG signals, making them less likely to generalize out of the box. Hence, the robustness, reliability, and interpretability of EEG-DL feature extractors remain open research problems.
@@ 2.3 Pillars of Foundation Modeling
In this section, we deconstruct the FM monolith into a set of universal, domain-agnostic building blocks or pillars. These pillars are minimally required to systematically compare diverse existing EEG-FMs and to build new ones.
@@
Representation of input data: EEG signals are acquired as a spatio-temporal data stream. Visual review of clinical EEGs is conducted primarily in this ‘native’ view. However, the computational interpretation and analysis of EEGs can benefit from alternative, perhaps more informative, representations of raw data. For example, representing EEGs as a temporal stream of short spatio-spectral data segments (10s each) could make it easier for EEG-FMs to learn latent frequency patterns and multi-scale information. Alternatively, representing EEGs using spatial covariance matrices can make it easier to extract latent connectivity patterns. Broadly, the choice of data representation at the input stage may greatly influence the quality and semantics of learned latent representations [neumann2021smart]. In biosignal domains where data are sampled from underlying physiological processes, constructing informative input data views may itself require significant offline transformations or even domain-specific modeling, such as inverse modeling [michel2019eeg]. Without these offline efforts, data-driven learning of latent representations via FMs may be infeasible.
@@
Architectural and functional design of FMs: At an architectural level, FMs are typically built from several non-linear transformer blocks that can be readily stacked to scale the model depth and size. Transformer blocks themselves contain multiple feed-forward layers followed by a multi-head attention module [vaswani2017attention]. Vision FMs may use convolutional blocks to exploit the desirable spatial inductive bias of convolutional operations [wu2021cvt]. Regardless of model size or low-level architectural choices, encoder-only FMs are understood to have only two high-level functional components: a backbone network and a final single fully connected layer. The backbone network serves as a feature extractor, providing low-dimensional embeddings (or representations) of the input data. The final layer, referred to as a task head, then uses these embeddings for task-specific purposes such as classification or regression.
@@
Self-supervised learning (SSL): In the SSL paradigm, supervision for training is derived strategically from unlabeled input data (pseudo-labels) rather than from human-curated labels, as in supervised learning [ericsson2022self]. SSL objectives (pretext tasks) are commonly either contrastive or generative. Contrastive pretext tasks use meaningfully augmented or noisy views of the input to learn robust data representations using encoder-only backbones [chen2020simple]. Generative pretext tasks mask portions of the input and reconstruct/generate the masked portions using encoder and decoder-based backbones [zhang2023survey]. In either case, SSL learns general intrinsic relationships within the data, thereby creating a task-agnostic backbone network that serves as a general feature extractor. The structure of representations learned by SSL tasks and general principles for effective SSL task design are under active study. Regardless, the value of SSL is particularly high in the EEG domain as it improves model generalizability [wagh2021domain] and partially alleviates the need for expert EEG annotations that are expensive to obtain, subjective, and error-prone [banville2021uncovering].
@@
Transfer learning and model adaptation: The backbone network enables the reuse, transfer, and adaptation of EEG features (i.e., data-driven EEG knowledge) across semantically overlapping tasks [thrun1998learning]. The proportion of the backbone adapted for a specific application is flexible. For example, the backbone may be entirely fixed/frozen (linear probing), partially updated, or entirely updated (fine-tuning) [kumar2022fine]. This adaptation process may use slower or layer-specific learning rates to preserve pretrained knowledge within the backbone. In either case, any pretext task-specific layers at the top are discarded and replaced with a new trainable linear head that performs classification or regression for the ‘downstream’ application. Using application-specific data splits and human-curated task-specific labels, the model is then trained further until convergence. The development of algorithms for robust and efficient adaptation is an active area of research. Broadly, transfer learning and model adaptation techniques obviate the need to train new EEG-DL models from scratch by leveraging knowledge from pre-existing models.
@@
Data and model scale: The scaling up of pretraining data volume, data diversity (i.e., the number of distinct data sources), and model size has contributed towards improved pretrained latent features, increased downstream task performance, sample efficiency, and model generalizability, among other properties [sun2017revisiting, dehghani2023scaling, kaplan2020scaling, yao2024towards, isik2024scaling, edwards2024scaling]. Empirical scaling laws have established power-law relationships between FM error rates and three key modeling levers, namely the pretraining data size, model size, and compute [kaplan2020scaling, rosenfeld2019constructive], which suggest that an exponential increase in data and model size can indeed improve FM performance, albeit with diminishing returns. Overall, the balance between data and model size to maximize performance under a fixed compute budget remains a critical consideration in FM research. It is noteworthy that vision and language FMs are often multi-billion-parameter models pretrained on an internet-scale collection of diverse datasets [oquab2023dinov2, wenzek2019ccnet, dosovitskiy2020image]. However, the relative scarcity of large-scale EEG corpora may limit scaling efforts in EEG-FM research.